Ankit is standing in the playing field at the back of his school. From where he is standing the angle of elevation of the bottom of the flagpole on the roof of the school is 35 degree. while the angle of elevation of the top of the flagpole is 40 degree. The distance from the ankit to the bottom of the flagpole is 37m and the distance from the ankit to the top of the flagpole is 40 m.
a) find the height of the flagpole
b0 How far along the level of the ground is ankit from the school

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

These are right triangles with the distance from Ankit to the flagpole being the hypotenuse of a right triangle, whose legs are the wall of the school and the distance from the school to him. So the angle of elevation of 35 degrees has the opposite side=height of school to bottom of flagpole, and the angle of elevation of 40 degrees has the opposite side=height of school to the top of the flagpole. The difference in these two is the height of the flagpole.
sin 35=x/37
x=37 sin 35=21.22 m, the height of the school
sin 40=x/40
x=40 sin 40=25.71 m
The flagpole is 4.49 m high.
The school is 21.22 m tall.
tan 35=21.22/y, where y is the distance of Ankit to the school
y=21.22/tan 35=30.3 m

Answer by MathTherapy(10551)


Ankit is standing in the playing field at the back of his school. From where he is standing the angle of elevation of the bottom of the flagpole on the roof of the school is 35 degree. while the angle of elevation of the top of the flagpole is 40 degree. The distance from the ankit to the bottom of the flagpole is 37m and the distance from the ankit to the top of the flagpole is 40 m.
a) find the height of the flagpole
b0 How far along the level of the ground is ankit from the school

Distance Ankit is from the school building: